How long does it take light to travel from the Sun to Venus? I know the distance is 1.03 x 10^8 and the speed of light is 300,000 kilometer per second.
d=rt
1.0 X 10^8 = (3.0 X 10^5)t
1.0 X 10^8 / 3.0 X 10^5
? X 10^3

Answer by Alan3354(69443)

How long does it take light to travel from the Sun to Venus? I know the distance is 1.03 x 10^8 and the speed of light is 300,000 kilometer per second.
d=rt
1.0 X 10^8 = (3.0 X 10^5)t
1.0 X 10^8 / 3.0 X 10^5
? X 10^3
----------
d = rt
t = d/r = 1.03E8/3E5
t = (1.03/3)*10^3 seconds
t = 1030/3 seconds
t = 343.333... seconds